OpenStudy (anonymous):

Find the height of a square pyramid with volume 243.73 ft3 and dimensions of base are 7.5 ft by 7.5 ft. A) 13 ft B) 8.67 ft C) 4.33 ft D) 9 ft

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

do you know the formula for the volume of a square based pyramid? :)

OpenStudy (anonymous):

V = a^2 h/3

OpenStudy (butterflydreamer):

yep! :) So in this case, "a" means the length of the base of the pyramid. So the question tells you a = 7.5 So since we know that you can plug it into the volume formula: V = a^2 * h/3 So V = (7.5)^2 *h/3 They also told you the volume was 243.73 so plug that into the formula aswell. 243.73 = (7.5)^2 * h/3

OpenStudy (butterflydreamer):

alright first. What does (7.5)^2 = ?

563blackghost (563blackghost):

Square Pyramid: \[V=a^2\frac{ h }{ 3 }\] Lets input what we know \[243.73ft^3=7.5^2\frac{ h }{ 3 }\] So we should work backwards so we start by simplifying the 7.5^squared after that you will divide the answer by 243.73 and get an irrational number which you will multiply by 3 to get the height

OpenStudy (anonymous):

h=12.9 ?

563blackghost (563blackghost):

Very Close!! its actually 13 xD

OpenStudy (butterflydreamer):

Well it'd be h = 12.998 So if we round it up it would be 13 :)

563blackghost (563blackghost):

Great Job!

OpenStudy (anonymous):

yes thats what i got thanks :)